# Barren-plateau free variational quantum simulation of $\mathbb{Z}_{2}$ lattice gauge theories

### v2

@@ III.2 Ground state estimation
In Fig. 6, we plot the optimization to find the ground state energy of the Hamiltonian in Eq. (4) with $J=5$, $m=1$, $\mu=2$ and charges placed at $(0,0)$ and $(0,2)$. This case is interesting as the corresponding TN calculation is stuck in a local minima, with this value plotted as a horizontal line. Thus the TN calculation for comparison requires the addition of penalty terms to the Hamiltonian, while this is not true for the VQE using either GI or ZZ ansatz. This can occur since TN approaches tend to be biased towards weakly correlated phases.
Figure 6: VQE optimization of the ground state. Here we consider the Hamiltonian in Eq. (4) with $J=5$, $m=1$, $\mu=2$ and charges placed at $(0,0)$ and $(2,0)$. The main curves show two instances of the full state-vector simulation for a single run of VQE optimized using SLSQP. The bold curve is a generic instance, while the translucent curve shows the case where the VQE escapes the local minima. The grey dashed line is the ground state energy obtained using TN methods and we can see the simulation is stuck in the same local minima. The black dashed line is the reference TN simulation with the addition of a penalty term $V=100$ that we compare our results with. Without the addition of penalty terms, the VQE is successful in converging to the ground state. In the inset we show results also for fixed numbers of shots $[500,1000,5000,10000]$ averaged over 1515 runs. In order to consider fixed numbers of shots, we optimize using SPSA. It seems the final energy does not depend on the number of shots. The black dashed line is the ground state obtained using tensor networks, while the blue dashed line is the average of the SLSQP runs. We note that on average, with this configuration SPSA attains values higher than if we optimized using SLSQP.
@@ III.3 String breaking
The results for string breaking are captured by the static potential $V(d)$, where $V(d)=E_{0}-E_{c}$, and $E_{c}$ is the ground state with charges in Fig. 7. The lattice spacing $d$ is taken with reference to the fermion at $(0,0)$, so $d$ counts the number of links between this and the other available fermion. This means that for $d=1,2,3$ there are two fermion sites and $d=4$ contains the furthest fermion from $(0,0)$ at $(3,1)$. The oscillations in $V(d)$ for the two fermions per fixed $d$ is due to the boundary conditions and superpositions of string configurations, where this effect is heightened at the critical $d=2$. Here the charge at $(2,0)$ still has the presence of a flux string, while for $(1,1)$ we find the string vanishes. The effects of left and right boundary conditions can be mitigated by considering longer ladder geometries, then studying the action of charges in the center. While for the superposition effects, where the flux string might take one of two paths, this choice can be imposed in the model.
@@
Fig. 7 also contains results obtained using IBM’s 156 qubit Heron r2 processor. We demonstrate the state preparation required for a single iteration of the VQE using 3 layers of the ZZ ansatz. Running the entire VQE would be too costly, so we consider the state preparation step that is required at every iteration. The ZZ ansatz was chosen as it is more hardware-efficient. Using Qiskit we perform the VQE up to some fixed number of iterations. The state we arrive at is not fully converged yet as this would require prolonged compute time, but is gauge-invariant, so we use this. The resultant 141 angles are then used to prepare the state on ibm_marrakesh, where the transpiled circuit has a depth of 525 and a $CZ$ count of 436. We use the default IBM error mitigation of zero-noise extrapolation (ZNE) and Twirled Readout Error eXtinction (TREX) [50], and take a plug-and-play approach to the experiments, where we choose the best of 5 runs. For the complicated circuits it is remarkable that the qualitative behavior of the VQE can be seen, but the results indicate that we should consider smaller system sizes. The initial angles for state preparation can be improved by using small penalty terms with the ZZ ansatz, since this would lead to faster convergence so reduced compute time. This would also introduce another aspect of the experiments that requires fine-tuning, that is not possible with the limited access to hardware.
@@ IV Discussion
One way to avoid barren plateaus is to use smart initialization strategies [22]; however a general strategy to identify these or to avoid local minima is not known. For LGTs on the other hand, we always initialize in some Gauss law sector and would know beforehand which to initialize in. Further, checking for Gauss law violations can alert as to whether the solution is a local minimum.
Another way is to constrain the searchable Hilbert space. This is done through the choice of ansatz, which is commonly designed to respect the symmetries of the problem. Solutions of LGTs are constrained to the gauge-invariant subspace of Hilbert space, and here we present an ansatz of gauge-invariant Hamiltonian terms that remains in this space. When the subspace spanned by the ansatz is polynomial in size, the problem can be classically simulated. All the while, the full space of gauge-invariant states is exponentially large [24, 25].
@@
Fundamentally, the VQE could never be performed in ideal conditions, so considering the effects of scaling with system size, number of shots and incorporating the effects of noise is vital for making such a determination. We find that even with low numbers of shots $\sim 500$, convergence to the ground state is close and the resulting state is gauge-invariant. We also find that when scaling with system size, this behavior remains constant, which indicates that for larger systems placed on quantum devices, the VQE could behave as we expect.
@@
In future work we would look to perform more experiments using quantum hardware. For smaller system sizes, it would be useful to perform some rounds of optimization in the VQE to see whether the resulting state remains gauge-invariant. This can be extended to observations of the string-breaking transition, where it would be interesting to see if this persists. We would also be interested to see the effect of small penalty terms on the cost landscape, since small terms in a region $\sim J$ can benefit the convergence of the $ZZ$ ansatz. Here one can also investigate the effect of small penalty terms in mitigating noise that may push the state out of the gauge-invariant subspace. We would also consider larger ladder lattices to mitigate the boundary effects. Some preliminary studies involving square lattices have also been performed, and show that we can extend this approach.
The behavior of the GI ansatz using current hardware versus the more hardware-efficient ZZ ansatz can also be analyzed, which would indicate the practicality of scaling such a VQE in the NISQ era. For those studies it would be beneficial to work more closely with the quantum hardware provider. It would also be beneficial to extend these studies to other platforms too. For instance the long-range multi-qubit interaction gates of the two ansatz may better suit ion-trap [51] or neutral atom based [52] quantum hardware. It is also interesting to explore applying the ZZ ansatz within the framework of measurement-based VQE [53], and explore the effectivity of having multi-qubit $Z$ rotation gates in the ansatz where the problem Hamiltonian has many-qubit terms.

### v3

@@ III.2 Ground state estimation
In Fig. 7, we plot the optimization to find the ground state energy of the Hamiltonian in Eq. (4) with $J=5$, $m=1$, $\mu=2$ and charges placed at $(0,0)$ and $(0,2)$. This case is interesting as the corresponding TN calculation is stuck in a local minima, with this value plotted as a horizontal line.
Figure 7: VQE optimization of the ground state. Here we consider the Hamiltonian in Eq. (4) with $J=5$, $m=1$, $\mu=2$ and charges placed at $(0,0)$ and $(2,0)$. The main curves show two instances of the full state-vector simulation for a single run of VQE optimized using SLSQP. The bold curve is a generic instance, while the translucent curve shows the case where the VQE escapes the local minima. The grey dashed line is the ground state energy obtained using TN methods and we can see the simulation is stuck in the same local minima. The black dashed line is the reference TN simulation with the addition of a penalty term $V=100$ that we compare our results with. Without the addition of penalty terms, the VQE is successful in converging to the ground state. In the inset we show results also for fixed numbers of shots $[500,1000,5000,10000]$ averaged over $15$ runs. In the software, rather optimizing with SPSA enables us to consider fixed numbers of shots. It seems the final energy does not depend on the number of shots. The black dashed line is the ground state obtained using tensor networks, while the blue dashed line is the average of the SLSQP runs. We note that on average, with this configuration SPSA attains values higher than if we optimized using SLSQP.
@@ III.3 String breaking
The results for string breaking are captured by the static potential $V(d)$, where $V(d)=E_{0}-E_{c}$, and $E_{c}$ is the ground state with charges in Fig. 8. The lattice distance $d$ is taken with reference to the fermion at $(0,0)$, so $d$ counts the number of links between this and the other available fermion. This means that for distances $d=1,2,3$ there are two possible fermion sites to place the static charge, whereas for $d=4$ there the only choice is site $(3,1)$. The different values in $V(d)$ for fixed distances $0<d<4$ are due to the different possible positions of the static charges and due to the corresponding boundary conditions, where this effect is heightened at the critical $d=2$. Here the charge at $(2,0)$ still has the presence of a flux string, while for $(1,1)$ we find the string vanishes. The effects of left and right boundary conditions can be mitigated by considering longer ladder geometries, then studying the action of charges in the center. While for the superposition effects, where the flux string might take one of two paths, this choice can be imposed in the model.
@@
Fig. 8 also contains results obtained using IBM’s 156 qubit Heron r2 processor. We demonstrate the state preparation required for a single iteration of the VQE using 3 layers of the ZZ ansatz. Running the entire VQE would be too costly, so we consider the state preparation step that is required at every iteration. The ZZ ansatz was chosen as it is more hardware-efficient. Using Qiskit we perform the VQE up to some fixed number of iterations. The state we arrive at is not fully converged yet as this would require prolonged compute time, but is gauge-invariant, so we use this. The resultant 141 angles are then used to prepare the state on ibm_marrakesh, where the transpiled circuit has a depth of 525 and a $CZ$ count of 436. We use the default IBM error mitigation of zero-noise extrapolation (ZNE) and Twirled Readout Error eXtinction (TREX) [vandenBerg2022Modelfree], and take a plug-and-play approach to the experiments, where we choose the best of 5 runs. For the complicated circuits it is remarkable that the qualitative behavior of the VQE can be seen, but the results indicate that we should consider smaller system sizes. The initial angles for state preparation can be improved by using small penalty terms with the ZZ ansatz, since this would lead to faster convergence so reduced compute time. This would also introduce another aspect of the experiments that requires fine-tuning, that is not possible with the limited access to hardware.
@@ IV Discussion
One way to avoid barren plateaus is to use smart initialization strategies [Puig2025]; however a general strategy to identify these or to avoid local minima is not known. For LGTs on the other hand, we always initialize in some Gauss law sector and would know beforehand which to initialize in. Further, checking for Gauss law violations can alert as to whether the solution is a local minimum.
Another way is to constrain the searchable Hilbert space. This is done through the choice of ansatz, which is commonly designed to respect the symmetries of the problem. Solutions of LGTs are constrained to the gauge-invariant subspace of Hilbert space, and here we present an ansatz of gauge-invariant Hamiltonian terms that remains in this space. When the subspace spanned by the ansatz is polynomial in size, the problem can be classically simulated. All the while, the full space of gauge-invariant states is exponentially large [Buividovich_2009, Tagliacozzo2014].
@@
Fundamentally, VQE would never be implemented in ideal conditions, so considering the effects of scaling with system size, number of shots and incorporating the effects of noise is vital for making such a determination. We find that even with low numbers of shots $\sim 500$, convergence to the ground state is close and the resulting state is gauge-invariant. We also find that when scaling with system size, this behavior remains constant, which indicates that for larger systems placed on quantum devices, the VQE could behave as we expect.
@@
In future work we would look to perform more experiments using quantum hardware. For smaller system sizes, it would be useful to perform some rounds of optimization in the VQE to see whether the resulting state remains gauge-invariant, and extending to observations of the string-breaking transition, where it would be interesting to see if this persists. We are also interested in the effects of small penalty terms on the cost landscape, since small terms in a region $\sim J$ can benefit the convergence of the $ZZ$ ansatz. Here one can also investigate the effect of small penalty terms in mitigating noise that may push the state out of the gauge-invariant subspace. Larger ladder lattice sizes would mitigate the boundary effects and we have performed preliminary studies involving square lattices, showing that we can extend this approach.
The behavior of the GI ansatz using current hardware versus the more hardware-efficient ZZ ansatz can also be analyzed, which would indicate the practicality of scaling such a VQE in the NISQ era. For those studies it would be beneficial to work more closely with the quantum hardware provider. It would also be beneficial to extend these studies to other platforms too. For instance the long-range multi-qubit interaction gates of the two ansatz may better suit ion-trap [Molmer2000] or neutral atom based [baker2021exploitinglongdistanceinteractionstolerating] quantum hardware. It is also interesting to explore applying the ZZ ansatz within the framework of measurement-based VQE [Ferguson2021VQE], and explore the effectivity of having multi-qubit $Z$ rotation gates in the ansatz where the problem Hamiltonian has many-qubit terms.